OpenStudy (anonymous):

sin x - sqrt3 cos x=1

OpenStudy (anonymous):

write as a single function of sine

OpenStudy (anonymous):

thank you but that is where I am getting confused. lol

OpenStudy (anonymous):

\[a\sin(x)+b\cos(x)=\sqrt{a^2+b^2}\sin(x+\theta)\] where \[\sin(\theta)=\frac{b}{\sqrt{a^2+b^2}}\] and \[\cos(\theta)=\frac{a}{\sqrt{a^2+b^2}}\]

OpenStudy (anonymous):

sometimes you can use \[\theta=\tan^{-1}(\frac{b}{a})\] but not always

OpenStudy (anonymous):

you good from there? this one is cooked up to be more or less easy

OpenStudy (anonymous):

not really these problems I tend to make ten times harder than it needs to be.

OpenStudy (anonymous):

in your case \(a=1,b=\sqrt3\) so \(\sqrt{a^2+b^2}=\sqrt{1+3}=\sqrt4=2\)

OpenStudy (anonymous):

actually \(b=-\sqrt3\) but it makes no difference when you square it

OpenStudy (anonymous):

as for \(\theta\), you have to find a number (angle if you prefer) where \[\sin(\theta)=\frac{b}{a}=-\frac{\sqrt3}{2}\] and \[\cos(\theta)=\frac{1}{2}\] this should look familiar

OpenStudy (anonymous):

oops i did not mean \(\sin(\theta)=\frac{b}{a}\) i meant \(\sin(\theta)=\frac{ b}{\sqrt{a^2+b^2}}\)

OpenStudy (anonymous):

so if the answer needs to be on an interval from 0 to 2pi i would add the period to the answer until i reach 2pi?

OpenStudy (anonymous):

we are not close to done yet what is \(\theta\)?

OpenStudy (anonymous):

oh lol they dont have give what theta is

OpenStudy (anonymous):

the instruction say give the answer in exact form on the interval of [0, 2pi) the question itself is sin theta - square root 3 cos theta = 1

OpenStudy (loser66):

you can square both sides to get \(sin^2 -2\sqrt{3} sin~ cos +3 cos^2 =1\) and separate to \(sin^2 + cos^2 +2 cos^2 -2 \sqrt{3}sin~ cos=1\), you have \(sin^2 + cos^2 =1\) cancel out with the right handside, therefore, just \(2cos^2 -2\sqrt{3}sin~cos=0\) factor 2cos out, \(2cos (cos -\sqrt{3}sin) =0\) so, cos =0 or \(cos -\sqrt{3}sin =0\) in that interval, cos =0 iff \(\theta =\dfrac{\pi}{2}~ or ~\theta=\dfrac{3\pi}{2}\) \(cos - \sqrt{3}sin =0\\cos = \sqrt{3}sin\\\dfrac{cos}{sin}= \sqrt{3}\\cot =\sqrt{3}\) you do the leftover

OpenStudy (anonymous):

thank you very much for your help